OpenStudy (anonymous):

-(1/4) sin (x/2) = 0 , [0,4pi] How do I find the value of x? Thank youu

hartnn (hartnn):

so, sin (x/2) =0 for which angles is sin =0 in the range 2 to 2 pi ??

hartnn (hartnn):

sorry, range, 0 to 2 pi***

OpenStudy (anonymous):

um.... I don't really know for the range... but the interval is the domain so @hartnn

hartnn (hartnn):

oh yes, ok. so for which angles is sin ratio =0 for domain of x from 0 to 2pi ?

OpenStudy (anonymous):

0 and pi

hartnn (hartnn):

and 2pi too, right ? so you know that sin 0 = 0, sin pi =0 and sin 2pi =0 ? that means, x/2 = 0 or x/2 = pi or x/2 = 2pi from these , what are the values of x we get ?

OpenStudy (anonymous):

oohhh I got 0, 2pi and 4pi is that right?

hartnn (hartnn):

absolutely! :)

OpenStudy (anonymous):

Thank you! :)

hartnn (hartnn):

welcome ^_^